Answer all the questions related to the exercises below. You can use Excel or work the exercise manually, as you prefer. You must detail the steps. That is, even with the support of the schedule, the process should be reflected in the worksheet

  1. 15% of students have a credit card balance higher than $5.500. Suppose you select 8 students to interview them about the use of your credit card (Binomial)
    1. What is the probability that 5 of the students will have a balance on their credit card higher than $5.500?
    2. That at least 3 have a balance on your credit card higher than $5.500?
Answer #1

Ans:

Use Binomial distribution with n=8 and p=0.15

P(x=k)=8Ck*0.15k*(1-0.15)8-k

a)

P(x=5)=8C5*0.155*0.853=0.0026

b)

P(at least 3)=P(x>=3)

=1-P(x<=2)

=1-P(x=0)-P(x=1)-P(x=2)

=1-8C0*0.150*0.858-8C1*0.151*0.857-8C2*0.152*0.856

=1-0.2725-0.3847-0.2376

=0.1052